OpenStudy (anonymous):

CALCULUS : Find the value of the Integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}x \sqrt{x^2-9}\]

OpenStudy (anonymous):

int x(x^2 -9)^ 1/2 1/3 (x^2-9)^3/2

OpenStudy (anonymous):

@ayoubEpst Please use the EQUATION EDITOR

OpenStudy (anonymous):

\[= 1/3 (x^2-9)\sqrt((x^2-9)) \] +c

OpenStudy (anonymous):

@ayoubEpst I have wolfram alpha too

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

I don't need the answer

OpenStudy (anonymous):

yes i can explain

OpenStudy (anonymous):

I hope you can

OpenStudy (anonymous):

\[\int\limits_{?}^{?} u'u^n\]

OpenStudy (anonymous):

you know the integral of that kind ??

OpenStudy (anonymous):

no i don't

OpenStudy (johnweldon1993):

I would just make a u-substitution here make u = x^2 - 9 so du = 2x dx and du/2 = xdx Now you have \[\large \frac{1}{2}\int \sqrt{u} du\]

OpenStudy (anonymous):

@ayoubEpst You're method is good but I just want to know all the fomula's from the table

OpenStudy (anonymous):

so you want that i give you the table ???

OpenStudy (anonymous):

Sorry for my post but the real question for this problem was to be solved by using the integration by parts method so....

OpenStudy (anonymous):

that's why I'm sutcked

OpenStudy (anonymous):

\[u'=x ; u=(x^2-9)^1/2\]

OpenStudy (anonymous):

@ayoubEpst I know all about the integration by parts method but I'm stucked by a point further

OpenStudy (anonymous):

@ayoubEpst continue the integration and you'll see what I'm talking about